OpenStudy (anonymous):

How would you factor 8x3+27

OpenStudy (mertsj):

Rewrite it like this: \[(2x)^3+(3)^3\] And then use this pattern: \[a^3+b^3=(a+b)(a^2-ab+b^2)\]

OpenStudy (anonymous):

do you mean \[8x ^{2}+3x+27\] or do you mean \[8x ^{3}+27\]

OpenStudy (anonymous):

How do you guys write like that >.< lol i can't figure that out!

OpenStudy (mertsj):

Use the equation button at the bottom of this box.

OpenStudy (tyteen4a03):

Alternatively learn LaTeX. :P

OpenStudy (anonymous):

How i would do the problem is i would automatically plug in the 8x, so it would be \[(8x \pm x) and (x \pm x)\]

OpenStudy (anonymous):

Oh then the equation i have to factor is \[8x ^{3}+27\]

OpenStudy (anonymous):

Actually im going to have to correct myself because i cant seem to figure out how i used to do this

OpenStudy (whpalmer4):

Jocelyn, good job using the equation editor - makes it much easier to read! Here's how I would factor it. \[8x^3+27\] has two cubes for coefficients. I'll take the cube roots as the coefficients of my first term: \[(8x^3+27)=(2x+3)(ax^2+bx+c)\] Now, let's multiply that out \[(8x^3+27)=2ax^3+2bx^2+2cx+3ax^2+3bx+3c\]\[=2ax^3+(2b+3a)x^2+(2c+3b)x+3c\] Let's figure out the values of a,b,c that make that fit our original. \[2a=8\]\[3c=27\]giving us\[a=4,c=9\]and \[8x^3+(2b+3(4))x^2+(2(9)+3b)x+27\] But we need those x^2 and x terms to vanish, so we solve \[2b+12=0\] giving us b=-6. Our factored equation is therefore \[(8x^3+27)=(2x+3)(4x^2-6x+9)\]

OpenStudy (whpalmer4):

You could also do polynomial division after choosing the first term. Harder to illustrate that :-)